Initially, a particle is moving at 4.50 m/s at an angle of 35.0° above the horizontal (+x axis). Two seconds later, its velocity is 5.79 m/s at an angle of 50.0° below the horizontal. What was the particle's average acceleration during these 2.00 seconds?

1. m/s2 (x component)

2. m/s2 (y component)
Answer #1

V1 = 4.5 cos35i + 4.5 sin35j = 3.69i + 2.58j

V2 = 5.79 cos50i - 5.79 sin50j = 3.72i - 4.44j

Average acceleration, a = (V2 - V1) /t

ax = 0.015 m/s^2

ay = - 3.51 m/s^2

Comment in case any doubt please rate my answer ....